# Rate-Distortion-Perception Tradeoff for Gaussian Vector Sources

### v1

@@ Abstract
This paper studies the rate-distortion-perception (RDP) tradeoff for a Gaussian vector source coding problem where the goal is to compress the multi-component source subject to distortion and perception constraints. The purpose of imposing a perception constraint is to ensure visually pleasing reconstructions. This paper studies this RDP setting with either the Kullback-Leibler (KL) divergence or Wasserstein-2 metric as the perception loss function, and shows that for Gaussian vector sources, jointly Gaussian reconstructions are optimal. We further demonstrate that the optimal tradeoff can be expressed as an optimization problem, which can be explicitly solved. An interesting property of the optimal solution is as follows. Without the perception constraint, the traditional reverse water-filling solution for characterizing the rate-distortion (RD) tradeoff of a Gaussian vector source states that the optimal rate allocated to each component depends on a constant, called the water-level. If the variance of a specific component is below the water-level, it is assigned a zero compression rate. However, with active distortion and perception constraints, we show that the optimal rates allocated to the different components are always positive. Moreover, the water-levels that determine the optimal rate allocation for different components are unequal. We further treat the special case of perceptually perfect reconstruction and study its RDP function in the high-distortion and low-distortion regimes to obtain insight to the structure of the optimal solution.
@@
Rate-distortion-perception function, lossy source coding, lossy compression, Gaussian vector sources, reverse water-filling
@@ I Introduction
The extension of classical rate-distortion (RD) theory to incorporate constraints on the distribution of the reconstruction samples has been studied in various works in the information theory literature; see e.g., [6] and references therein. More recently, Theis and Wagner [7] present a one-shot coding theorem by means of the strong functional representation lemma (SFRL) [8] to establish the operational validity of the RDP function [5]. In [9], the authors establish analytic properties of the RDP function for the special case of (scalar) Gaussian sources, with a quadratic distortion function and a perception loss function of either Kullback–Leibler (KL) divergence or Wasserstein-2 distance between the source and the reconstruction distributions. The role of common randomness in the study of RDP function has been studied in [10, 11]. Furthermore, the distortion-perception tradeoff with a squared error distortion and Wasserstein-2 perception loss, but without an explicit compression rate constraint, has been studied in [12, 13], where it is shown that the entire tradeoff curve can be achieved by interpolating the two extremal reconstructions based on a given representation. Other related works include [14, 15].
@@
This paper studies the RDP function of a Gaussian vector source under a squared error distortion and either KL divergence or Wasserstein-2 distance as the perception loss metric. Our result is thus an extension of prior work [9] on scalar Gaussian sources to the case of vector sources. We start by demonstrating the optimality of jointly Gaussian reconstructions for Gaussian vector sources in the RDP setting. We then show that by decomposing the Gaussian vector source using the unitary transformation obtained from the eigenvalue decomposition of its covariance matrix, it is possible to derive an achievable RDP function of the Gaussian vector source in term of the RDP functions of its constituent scalar components. The optimality of this achievable scheme can be established by a converse proof. This means that the characterization of the optimal RDP function can be formulated as an optimization problem. We explicitly derive the solution of the optimization problem and investigate structural properties of the optimal solution.
@@
The optimal RDP function for the Gaussian vector source has the following interesting property. Without the perception constraint, the rate-distortion function of a parallel Gaussian source model has a classical reverse water-filling characterization [16, Thm 10.3], where the optimal rate allocation across the components is computed according to a distortion dependent parameter called water-level. A positive rate is assigned to those components that have a variance above this parameter. Any component whose variance is below the water-level has a zero rate; see Fig. 1(a). However, with a perception constraint, we observe a qualitatively different solution as shown in Fig. 1(b). First, unlike the case of reverse water-filling, the associated water-level for each component can be different and is characterized as a solution to a set of equations. Second, while reverse water-filling assigns zero rate to those source components whose variances are below the water-level, all components in the RDP setting are assigned a non-zero rate as long as both the distortion and perception constraints are active.
@@
Figure 1: (a) Without a perception constraint, the traditional reverse water-filling solution for a parallel Gaussian source fixes a constant water-level. When the variance of a specific component is less than the water-level, it is assigned zero rate. (b) With an active perception constraint, unequal water-levels are assigned to different components. The variance of each component is always greater than the corresponding water-level. Every component has a positive rate.
@@ III Traditional Reverse Water-Filling
The classical rate-distortion theory for a parallel Gaussian source states that the optimal rate allocated to each component depends on a constant parameter, called water-level, as shown in Fig. 1(a). The water-level also represents the distortion allowed at those components whose variances are above the water-level. For a given distortion $D$, let $\nu(D)$ be the solution to the equation
@@
To simplify notation, we can redefine the water-level as $\gamma_{\ell}^{*}(D,\infty)$ in order to account for the components whose variances are below the water-level. If $\lambda_{\ell}$ is below $\nu(D)$ for some $\ell$, then we set $\gamma_{\ell}^{*}(D,\infty)=\lambda_{\ell}$ and assign zero rate to this component. Two special cases of the above theorem are of special interest.
@@
In the high-distortion regime, we have that for sufficiently small $\epsilon>0$
@@
where $\lambda^{\max}=\max_{\ell}\lambda_{\ell}$. Let $L^{\max}$ denote the set of indices where their corresponding eigenvalues are equal to $\lambda^{\max}$. Then, the water-levels are given by
$\gamma_{\ell}^{*}\left(\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,\infty\right)$ $=$ $\lambda_{\ell},\qquad\forall\ell\in\{1,\ldots,L\}\backslash L^{\max},$ (23a)
$\gamma^{*}_{\ell^{\max}}\left(\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,\infty\right)$ $=$ $\lambda^{\max}-\frac{\epsilon}{|L^{\max}|},\;\;\forall\ell^{\max}\in L^{\max}.$ (23b)
@@
See Appendix A-1. ∎
The above proposition states that in the high-distortion compression, a positive rate is only assigned to the components with the largest eigenvalue.
@@
In the low-distortion regime, we have that for a sufficiently small $\epsilon>0$
@@
where the water-levels are given by
@@
See Appendix A-2. ∎
For low-distortion compression, according to the above proposition, the same water-level is assigned to all components.
@@ IV-A Optimality of Gaussian Reconstruction
We first present a result indicating that for the two perception metrics (15) and (16) considered in this paper and for a Gaussian vector source, jointly Gaussian reconstruction is optimal.
@@ IV-B RDP Function with KL Divergence as Perception Metric
The main idea is to construct a new Gaussian random vector $\hat{Z}$ and to use the channel simulation result of [8] to communicate $\hat{Z}$ to the decoder at a rate of $R$. The new random vector $\hat{Z}$ is designed to be correlated with $Z$ in a very specific way in order to satisfy the distortion and perception constraints $D$ and $P$, respectively. The correlation between $Z$ and $\hat{Z}$ is controlled by two sets of parameters, $\{\gamma_{\ell}\}_{\ell=1}^{L}$ and $\{\hat{\lambda}_{\ell}\}_{\ell=1}^{L}$, such that $0<\gamma_{\ell}\leq\lambda_{\ell}$ and $0<\hat{\lambda}_{\ell}\leq\lambda_{\ell}$. The optimal values of these parameters will be determined later.
@@
Now, to derive the one-shot RDP function $R^{o}(D,P)$, we can make use a consequence of the SFRL [8, Theorem 1] to show that when common randomness $K$ is available at both the encoder and decoder, there exists a channel simulation scheme that allows $\hat{Z}_{\ell}$ to be reconstructed at the decoder at a communication rate of
@@
${R}_{\ell}(\gamma_{\ell})$ $=$ $\frac{1}{2}\log\left(\frac{\lambda_{\ell}}{\gamma_{\ell}}\right)+\log\left(\frac{1}{2}\log\left(\frac{\lambda_{\ell}}{\gamma_{\ell}}\right)+1\right)+5,$ (31)
@@
For the asymptotic setting, the achievable scheme is identical, except that we compress a block of $n$ samples together. As $n\rightarrow\infty$, the logarithm and the constant terms in  (31) can be neglected. This leads to an upper bound for $R^{\infty}(D,P)$, which is equal to $R(D,P)$. This upper bound turns out to be tight, i.e., a converse can be proved. This gives the following characterization of $R(D,P)$.
@@
s.t. $0<\gamma_{\ell}\leq\lambda_{\ell},$ (34e)
$0\leq\hat{\lambda}_{\ell},$
@@
An interpretation of the above is as follows. For a given $(D,P)$, let $\gamma_{\ell}^{*}(D,P)$ and $\hat{\lambda}^{*}_{\ell}(D,P)$, $\ell\in\{1,\ldots,L\}$, be the optimal solution to (34). Comparing this with (21), it can be seen that $\gamma^{*}_{\ell}(D,P)$ can be interpreted as the water-level for the $\ell$-th component, which determines the rate allocated to that component according to (34a); see Fig. 1(b).
@@ IV-C Generalized Water-filling with KL Divergence as Perception Metric
$\nu_{1}\left(-\sqrt{\frac{\lambda_{{\ell}}-\gamma^{*}_{{\ell}}(D,P)}{\hat{\lambda}^{*}_{{\ell}}(D,P)}}+1\right)+\frac{\nu_{2}}{2}\left(\frac{1}{\lambda_{{\ell}}}-\frac{1}{\hat{\lambda}^{*}_{{\ell}}(D,P)}\right)-\eta_{\ell}=0.$ (36)
@@
$\gamma^{*}_{\ell}(D,P)$ $=$ $\frac{-2\lambda_{\ell}\nu_{1}(1+2\lambda_{\ell}\nu_{1})-\nu_{2}+\sqrt{(\nu_{2}+2\lambda_{\ell}\nu_{1}+4\lambda_{\ell}^{2}\nu_{1}^{2})^{2}+16\lambda_{\ell}^{2}\nu_{1}^{2}(\nu_{2}+2\lambda_{\ell}\nu_{1})(\nu_{2}-1)}}{8\lambda_{\ell}\nu_{1}^{2}(-1+\nu_{2})}.$
@@
The expressions (IV-C) and (37) give us the following generalized reverse water-filling interpretation of the optimal RDP solution. At given distortion constraint $D$ and perception constraint $P$, each component of the source with variance $\lambda_{\ell}$ is reconstructed by $\hat{Z}_{\ell}$ having a variance $\hat{\lambda}^{*}_{\ell}(D,P)$. Because $\gamma^{*}_{\ell}(D,P)$ is the variance of the MMSE estimate of $Z_{\ell}$ given $\hat{Z}_{\ell}$, this requires a rate of $\frac{1}{2}\log\left(\frac{\lambda_{\ell}}{\gamma^{*}_{\ell}(D,P)}\right)$. The parameters $\hat{\lambda}^{*}_{\ell}(D,P)$ and $\gamma^{*}_{\ell}(D,P)$ are chosen to satisfy the distortion and perception constraints. As already mentioned, $\gamma^{*}_{\ell}(D,P)$ can be thought of as the water-level, cf. (21).
@@
so every component of the source is always allocated a non-zero rate regardless of the distortion constraint—unlike the traditional reverse water-filling solution, where a component may be allocated zero rate if its variance is below the water-level. Moreover, in contrast to the traditional reverse water-filling, the distortion of each component (i.e., $D_{\ell}(\gamma_{\ell}^{*}(D,P),\hat{\lambda}_{\ell}^{*}(D,P))$) may not be the same across the different components. So, an unequal-distortion allocation may be optimal when both perception and distortion constraints are active.
It is also possible that either the distortion or the perception constraint is not active. If the distortion constraint is active while the perception constraint is inactive, i.e., $\nu_{1}>0$ and $\nu_{2}=0$, and $\eta_{\ell}=\eta^{\prime}_{\ell}=0$ for all $\ell\in\{1,\ldots,L\}$, then (35) and (36) yield the traditional reverse water-filling solution. Specifically, the water-level is given by $\min\{\frac{1}{2\nu_{1}},\lambda_{\ell}\}$ where $\nu_{1}$ satisfies the following:
@@
An interesting observation is that based on (41) and (43), we see that when the perception constraint is active, it is either that all the components are allocated positive rate, or that all the components are allocated zero rate. This means that the situation in the traditional reverse water-filling, where some of the water-levels are below the eigenvalues while others are equal to the eigenvalues, cannot happen, when the perception constraint is active.
@@
Let $(D,P)$ be a given distortion and perception constraints that are strictly feasible. The optimal solution of (34) with KL divergence as the perception metric is given as follows:
@@
    $\left\{\hat{\lambda}_{\ell}\ \left|\ \sum_{\ell=1}^{L}{P}_{\ell}(\hat{\lambda}_{\ell})\leq P,\ \ \sum_{\ell=1}^{L}\lambda_{\ell}+\hat{\lambda}_{\ell}\leq D,\ \ \hat{\lambda}_{\ell}\geq 0\right.\right\}.$ (47)

### v2

@@ Abstract
This paper studies the rate-distortion-perception (RDP) tradeoff for a Gaussian vector source coding problem where the goal is to compress the multi-component source subject to distortion and perception constraints. Specifically, the RDP setting with either the Kullback-Leibler (KL) divergence or Wasserstein-2 metric as the perception loss function is examined, and it is shown that for Gaussian vector sources, jointly Gaussian reconstructions are optimal. We further demonstrate that the optimal tradeoff can be expressed as an optimization problem, which can be explicitly solved. An interesting property of the optimal solution is as follows. Without the perception constraint, the traditional reverse water-filling solution for characterizing the rate-distortion (RD) tradeoff of a Gaussian vector source states that the optimal rate allocated to each component depends on a constant, called the water level. If the variance of a specific component is below the water level, it is assigned a zero compression rate. However, with active distortion and perception constraints, we show that the optimal rates allocated to the different components are always positive. Moreover, the water levels that determine the optimal rate allocation for different components are unequal. We further treat the special case of perceptually perfect reconstruction and study its RDP function in the high-distortion and low-distortion regimes to obtain insight to the structure of the optimal solution.
@@
Rate-distortion-perception function, lossy source coding, lossy compression, Gaussian vector sources, reverse water-filling.
@@ I Introduction
The extension of classical rate-distortion (RD) theory to incorporate constraints on the distribution of the reconstruction samples has been studied in various works in the information theory literature; see e.g., [6] and references therein. More recently, Theis and Wagner [7] present a one-shot coding theorem by means of the strong functional representation lemma (SFRL) [8] to establish the operational validity of the RDP function [5]. In [9], the authors establish analytic properties of the RDP function for the special case of (scalar) Gaussian sources, with a quadratic distortion function and a perception loss function of either Kullback–Leibler (KL) divergence or Wasserstein-2 distance between the source and the reconstruction distributions. The role of common randomness in the study of RDP function has been reported in [10, 11]. Furthermore, the distortion-perception tradeoff with a squared error distortion and Wasserstein-2 perception loss has been studied in [12, 13], where it is shown that the entire tradeoff curve can be achieved by interpolating the two extremal reconstructions based on a given representation. Other related works include [14, 15].
@@
This paper studies the RDP function of a Gaussian vector source under a squared error distortion and either KL divergence or Wasserstein-2 distance as the perception loss metric. Our result is thus an extension of prior work [9] on scalar Gaussian sources to the case of vector sources. We start by demonstrating the optimality of jointly Gaussian reconstructions for Gaussian vector sources in the RDP setting. We then show that by decomposing the Gaussian vector source using the unitary transformation obtained from the eigenvalue decomposition of its covariance matrix, it is possible to derive an achievable RDP function of the Gaussian vector source in term of the RDP functions of its constituent scalar components. The optimality of this achievable scheme can be established by a converse proof. This means the characterization of the optimal RDP function can be formulated as an optimization problem. We explicitly derive the solution of the optimization problem and investigate structural properties of the optimal solution.
@@
The optimal RDP function for the Gaussian vector source has the following interesting property. Without the perception constraint, the rate-distortion function of a parallel Gaussian source model has a classical reverse water-filling characterization [16, Thm 10.3], where the optimal rate allocation across the components is computed according to a distortion dependent parameter called water level. A positive rate is assigned to those components that have a variance above this parameter. Any component whose variance is below the water level has a zero rate; see Fig. 1(a). However, with a perception constraint, we observe a qualitatively different solution as shown in Fig. 1(b). First, unlike the case of reverse water-filling, the associated water level for each component can be different and is characterized as a solution to a set of equations. Second, while reverse water-filling assigns zero rate to those source components whose variances are below the water level, all components in the RDP setting are assigned a non-zero rate as long as both distortion and perception constraints are active.
@@
Figure 1: (a) Without a perception constraint, the traditional reverse water-filling solution for a parallel Gaussian source fixes a constant water level. When the variance of a specific component is less than the water level, it is assigned zero rate. (b) With an active perception constraint, unequal water levels are assigned to different components. The variance of each component is always greater than the corresponding water level. Every component has a positive rate.
@@ III Traditional Reverse Water-Filling
The classical rate-distortion theory for a parallel Gaussian source states that the optimal rate allocated to each component depends on a constant parameter, called water level, as shown in Fig. 1(a). The water level also represents the distortion allowed at those components whose variances are above the water level. For a given distortion $D$, let $\nu(D)$ be the solution to the equation
@@
To simplify notation, we can redefine the water level as $\gamma_{\ell}^{*}(D,\infty)$ in order to account for the components whose variances are below the water level. If $\lambda_{\ell}$ is below $\nu(D)$ for some $\ell$, then we set $\gamma_{\ell}^{*}(D,\infty)=\lambda_{\ell}$ and assign zero rate to this component. Two special cases of the above theorem are of particular interest.
@@
In the high-distortion regime, we have that for sufficiently small $\epsilon>0$,
@@
where $\lambda^{\max}=\max_{\ell}\lambda_{\ell}$. Let $\mathcal{L}^{\max}$ denote the set of indices where their corresponding eigenvalues are equal to $\lambda^{\max}$. Then, the water levels are given by
$\gamma_{\ell}^{*}\left(\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,\infty\right)$ $=$ $\lambda_{\ell},\;\;\forall\ell\in\{1,\ldots,L\}\backslash\mathcal{L}^{\max},$ (23a)
$\gamma^{*}_{\ell^{\max}}\left(\sum_{\ell=1}^{L}\lambda_{\ell}-\epsilon,\infty\right)$ $=$ $\lambda^{\max}-\frac{\epsilon}{|\mathcal{L}^{\max}|},\forall\ell^{\max}\in\mathcal{L}^{\max}.$
@@
See Appendix A-A. ∎
The above proposition states that in high-distortion compression, a positive rate is only assigned to the components with the largest eigenvalue.
@@
In the low-distortion regime, we have that for a sufficiently small $\epsilon>0$,
@@
where the water levels are given by
@@
See Appendix A-B. ∎
For low-distortion compression, according to the above proposition, the same water level is assigned to all components.
@@ IV-A Optimality of Gaussian Reconstruction
We first present a result showing that for the two perception metrics (15) and (16) and for a Gaussian vector source, jointly Gaussian reconstruction is optimal.
@@ IV-B RDP Function with KL Divergence as Perception Metric
The main idea is to construct a new Gaussian random vector $\hat{Z}$ and to use the channel simulation result of [8] to communicate $\hat{Z}$ to the decoder at a rate of $R$. The new random vector $\hat{Z}$ is designed to be correlated with $Z$ in a very specific way in order to satisfy the distortion and perception constraints $D$ and $P$, respectively. The correlation between $Z$ and $\hat{Z}$ is controlled by two sets of parameters, $\{\gamma_{\ell}\}_{\ell=1}^{L}$ and $\{\hat{\lambda}_{\ell}\}_{\ell=1}^{L}$, such that $0<\gamma_{\ell}\leq\lambda_{\ell}$ and $0<\hat{\lambda}_{\ell}\leq\lambda_{\ell}$. The optimal values of these parameters are determined later.
@@
Now, to derive an upper bound on the one-shot RDP function $R^{o}(D,P)$, we can use a consequence of the SFRL [8, Theorem 1] to show that when common randomness $K$ is available at both the encoder and decoder, there exists a channel simulation scheme that allows $\hat{Z}_{\ell}$ to be reconstructed at the decoder at a communication rate of
@@
${R}_{\ell}(\gamma_{\ell})$ $=$ $\frac{1}{2}\log\left(\frac{\lambda_{\ell}}{\gamma_{\ell}}\right)+\log\left(\frac{1}{2}\log\left(\frac{\lambda_{\ell}}{\gamma_{\ell}}\right)+1\right)+5,$
@@
For the asymptotic setting, the achievable scheme is identical, except that we compress a block of $n$ samples together. As $n\rightarrow\infty$, the logarithm and the constant terms in  (IV-B) can be neglected. This leads to an upper bound for $R^{\infty}(D,P)$, which is equal to $R(D,P)$. This upper bound turns out to be tight, i.e., a converse can be proved. This gives the following characterization of $R(D,P)$.
@@
s.t. $0<\gamma_{\ell}\leq\lambda_{\ell}$ (36e)
$0\leq\hat{\lambda}_{\ell}$
@@
An interpretation of the above is as follows. For a given $(D,P)$, let $\gamma_{\ell}^{*}(D,P)$ and $\hat{\lambda}^{*}_{\ell}(D,P)$, $\ell\in\{1,\ldots,L\}$, be the optimal solution to (36). Comparing this with (21), it can be seen that $\gamma^{*}_{\ell}(D,P)$ can be interpreted as the water level for the $\ell$-th component, which determines the rate allocated to that component according to (36a); see Fig. 1(b).
@@ IV-C Generalized Water-filling with KL Divergence as Perception Metric
$\nu_{1}\left(-\sqrt{\frac{\lambda_{{\ell}}-\gamma^{*}_{{\ell}}(D,P)}{\hat{\lambda}^{*}_{{\ell}}(D,P)}}+1\right)+$
$\hskip 56.9055pt\frac{\nu_{2}}{2}\left(\frac{1}{\lambda_{{\ell}}}-\frac{1}{\hat{\lambda}^{*}_{{\ell}}(D,P)}\right)-\eta_{\ell}=0.$ (38)
@@
$\gamma^{*}_{\ell}(D,P)=\frac{-2\lambda_{\ell}\nu_{1}(1+2\lambda_{\ell}\nu_{1})-\nu_{2}+\sqrt{\Delta(\nu_{1},\nu_{2})}}{8\lambda_{\ell}\nu_{1}^{2}(-1+\nu_{2})},$
where
$\Delta(\nu_{1},\nu_{2})$ $=(\nu_{2}+2\lambda_{\ell}\nu_{1}+4\lambda_{\ell}^{2}\nu_{1}^{2})^{2}$
$\hskip 14.22636pt+16\lambda_{\ell}^{2}\nu_{1}^{2}(\nu_{2}+2\lambda_{\ell}\nu_{1})(\nu_{2}-1).$ (42)
@@
The expressions (IV-C) and (39) give us the following generalized reverse water-filling interpretation of the optimal RDP solution. At given distortion constraint $D$ and perception constraint $P$, each component of the source with variance $\lambda_{\ell}$ is reconstructed by $\hat{Z}_{\ell}$ having a variance $\hat{\lambda}^{*}_{\ell}(D,P)$. Because $\gamma^{*}_{\ell}(D,P)$ is the variance of the MMSE estimate of $Z_{\ell}$ given $\hat{Z}_{\ell}$, this requires a rate of $\frac{1}{2}\log\left(\frac{\lambda_{\ell}}{\gamma^{*}_{\ell}(D,P)}\right)$. The parameters $\hat{\lambda}^{*}_{\ell}(D,P)$ and $\gamma^{*}_{\ell}(D,P)$ are chosen to satisfy the distortion and perception constraints. As already mentioned, $\gamma^{*}_{\ell}(D,P)$ can be thought of as the water level, cf. (21).
@@
so every component of the source is always allocated a non-zero rate regardless of the distortion constraint—unlike the traditional reverse water-filling solution, where a component may be allocated zero rate if its variance is below the water level. Moreover, in contrast to the traditional reverse water-filling solution, the distortion of each component (i.e., $D_{\ell}(\gamma_{\ell}^{*}(D,P),\hat{\lambda}_{\ell}^{*}(D,P))$) may not be the same across the different components. So, the optimal distortion allocation across the components may be unequal when both perception and distortion constraints are active.
It is also possible that either the distortion or the perception constraint is not active. If the distortion constraint is active while the perception constraint is inactive, i.e., $\nu_{1}>0$ and $\nu_{2}=0$, and $\eta_{\ell}=\eta^{\prime}_{\ell}=0$ for all $\ell\in\{1,\ldots,L\}$, then (37) and (38) yield the traditional reverse water-filling solution. Specifically, the water level is given by $\min\{\frac{1}{2\nu_{1}},\lambda_{\ell}\}$ where $\nu_{1}$ satisfies the following:
@@
An interesting observation is that based on (44) and (46), we see that when the perception constraint is active, it is either that all the components are allocated positive rate, or that all the components are allocated zero rate. This means that the situation in the traditional reverse water-filling, where some of the water levels are below the eigenvalues while others are equal to the eigenvalues, cannot happen, when the perception constraint is active.
@@
Let $(D,P)$ be strictly feasible distortion and perception constraints. The optimal solution of (36) with KL divergence as the perception metric is given as follows:
@@
    $\Bigg{\{}\hat{\lambda}_{\ell}\Bigg{|}\sum_{\ell=1}^{L}{P}_{\ell}(\hat{\lambda}_{\ell})\leq P,\sum_{\ell=1}^{L}\lambda_{\ell}+\hat{\lambda}_{\ell}\leq D,\hat{\lambda}_{\ell}\geq 0\Bigg{\}}.$ (50)